OpenStudy (anonymous):

The probability that Jacqueline will be elected to the students’ council is 0.6, and the probability that she will be selected to represent her school in a public-speaking contest is 0.75. What is the probability that Jacqueline is either elected to the students’ council or picked for the public-speaking contest?

OpenStudy (anonymous):

was thinking: 0.6*0.25 + 0.75*0.4 = 0.45 but the answer is 0.85

OpenStudy (anonymous):

assuming they are independent (big assumption) you take \[.6+.75-.6\times .75\]

OpenStudy (anonymous):

which is not .75

OpenStudy (anonymous):

i mean it is not .85

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

it is in fact .9

OpenStudy (anonymous):

i do know they are independant tho

OpenStudy (anonymous):

well fine. then answer is .9

OpenStudy (anonymous):

if they are independent then \[P(A\cap B)=P(A)P(B)\]

OpenStudy (anonymous):

P( A intersection B) was given at 0.5 the previous question

OpenStudy (anonymous):

??

OpenStudy (anonymous):

in that case they are not independent

OpenStudy (anonymous):

The probability that Jacqueline will be elected to the students’ council is 0.6, and the probability that she will be selected to represent her school in a public-speaking contest is 0.75. The probability of Jacqueline achieving both of these goals is 0.5. Sorry i didnt type some of the question

OpenStudy (anonymous):

because if they were independent then the probability of the intersection is .6*.75=.45

OpenStudy (anonymous):

ooooooooooooooooooh

OpenStudy (anonymous):

yes so they are independant

OpenStudy (anonymous):

no no no no sorry no

OpenStudy (anonymous):

IF they were independent then the probability of the intersection would be .45 but it is .5

OpenStudy (anonymous):

what you are supposed to do in this case is say .6+.75-.5=.85

OpenStudy (anonymous):

oh so they are dependant?

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

omg wow

OpenStudy (anonymous):

easy to check. if they were independent then the probability of the intersection would be the product of the probabilities. but is is not

OpenStudy (anonymous):

.6*.75=.45 not .5

OpenStudy (anonymous):

yes so they are not mutually exclusive

OpenStudy (anonymous):

be that as it may all you have to do is say \[P(A)+P(B)-P(A\cap B)=.6+.75-.5=.85\]

OpenStudy (anonymous):

well those are separate things.

OpenStudy (anonymous):

if they were mutually exclusive then \[P(A\cap B)=0\]

OpenStudy (anonymous):

so no they are not mutually exclusive, nor are they independent. they are dependent

OpenStudy (anonymous):

ahh yes thx makes soo much more sense now

OpenStudy (anonymous):

yw